i need to know the solution set for this rational equation
4/x=16/5x+4
thanks

Answer by stanbon(75887)

i need to know the solution set for this rational equation
4/x=16/5x+4
---
Cross-multiply to get:
20x+16 = 16x
4x = -16
x = -4
----------------
Cheers,
Stan H.